In a triangle ABC , If 'O' is the circumcentre and 'H' is the orthocentre,ย  then show thatย 
1) OA + OB + OC = OHย 
& 2) HA+HB+HC =2HO.
๐Ÿค”๐Ÿค”๐Ÿค”

Grade:11

1 Answers

Arun
5 years ago
we know that O'G =2GOย  where G is the centroid of trianglelet a point D between B and CSo, OD =(OB+OC )/2now OA + OB + OC = OA + 2ODwe know that G devide the point A and mib point of apposite side(D) in ratiioย  2:1So, OG =(OA 2OD)/(2+1)So, OA + OB + OC =3OG=2OG + OG=GO'+OG =OO'